OpenStudy (hijacktrolo):

Which relationship is always true for the angles r, x, y, and z of triangle ABC? x + z = y 180 degrees − x = r x + y + z = 180 degrees x + y + z = 90 degrees please help

OpenStudy (hijacktrolo):

x + z = y 180 degrees − x = r x + y + z = 180 degrees x + y + z = 90 degrees